OpenStudy (anonymous):

f:x=4x+m and f^-1:x=nx+3/4. Find the values of m and n. @ganeshie8

OpenStudy (anonymous):

f:x means function of x?

OpenStudy (mathmath333):

is the answer -3 and 1/4

ganeshie8 (ganeshie8):

yes

OpenStudy (mathmath333):

find the inverse of \(f(x)=4x+m\)

OpenStudy (anonymous):

y=4x+m

OpenStudy (anonymous):

\[\frac{ y-m }{ 4 }=x\]

OpenStudy (anonymous):

\[f ^{-1}(x)=\frac{ x-m }{ 4 }\]

OpenStudy (mathmath333):

ok compare the two inverse equations now

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{nx+\dfrac{3}{4}=\dfrac{x-m}{4} \hspace{.33em}\\~\\ \implies nx+\dfrac{3}{4}=x\cdot\dfrac{1}{4}+\dfrac{-m}{4} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }=\frac{ -m }{ 4 }\]

OpenStudy (anonymous):

\[-4m=12\]

OpenStudy (anonymous):

\[m=-3\]

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{nx+\dfrac{3}{4}=\dfrac{x-m}{4} \hspace{.33em}\\~\\ \implies \color{red}{n}x+\color{blue}{\dfrac{3}{4}}=x\cdot\color{red}{\dfrac{1}{4}}+\color{blue}{\dfrac{-m}{4}} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Thnx @mathmath333 :)

OpenStudy (mathmath333):

yw